# Alessandro Galeazzi1, Pujan Paudel2, Mauro Conti1, Emiliano De Cristofaro3, Gianluca Stringhini2

### v3

@@ I Introduction
Beyond anecdotal evidence, researchers have begun looking into shadow banning through technical and statistical analyses (see Section II). Overall, prior work suggests that shadow banning is relatively widespread across social media platforms [29, 7]. However, the lack of transparency in applying these techniques raises serious concerns. This prompts the need for systematic techniques to detect and measure visibility alterations and assess whether platforms employ shadow banning to favor specific narratives, limit dissenting voices, or prioritize corporate interests.
@@
In this paper, we aim to address this gap and evaluate the presence of visibility alterations in the wild using view counts, a largely underexplored metric in related work introduced on Twitter in 2022 and once available through the now mostly defunct API. As academic access to the API has been restricted since June 2023 [34, 22], we turn to two publicly available datasets that include view counts: 1) almost 35M tweets related to the 2024 US Presidential Election, released by Balasubramanian et al. [4], and 2) over 17M tweets related to military aid in the Ukraine-Russia war, released by Baqir et al. [5]. [^1] As Twitter rebranded to X in July 2023, tweets are now called posts. Since one of our datasets was collected before the rebranding and one after, to ease presentation, in the rest of the paper, we stick to “tweets.” Our methodology relies on data analysis and network science techniques to uncover traces of shadow banning at the content, user, and network levels, shedding light on visibility alterations. Although the inaccessibility of the platform’s recommendation algorithm prevents us from establishing an explicit causal link to algorithmic design, our findings provide reasonable evidence in support of this hypothesis. While alternative explanations are possible, we discuss why they are unlikely to hold in Section VI.
@@ II Related Work
Specifically related to shadow banning, Le Merrer et al. [29] investigate different algorithmic interventions on Twitter aimed to reduce the visibility of certain content or accounts without outright removal. Their findings highlight the widespread presence of traces of visibility interventions across various user groups and communities. Compared to our work, their study focuses on account-level interventions based on appearances in suggestions, search results, or discussion cascades, without considering alterations at the content level.
@@
III Dataset
2024 US Elections. The second dataset was collected by Balasubramanian et al. [4] using a custom-built scraper. [^3] The dataset is available at https://github.com/sinking8/usc-x-24-us-election. The authors curated a list of 44 keywords related to the 2024 US Elections, including ‘2024 Elections,’ ‘Donald Trump,’ ‘letsgobrandon,’ ‘ultramaga,’ etc. Following retrieval, we filter the data as discussed in Section IV-A and obtain approximately 35 million tweets with valid view counts by 3.8 million users.
As we apply several transformations to the datasets throughout the different stages of our analysis (see Section IV), we provide summary statistics for the datasets used in the main steps of this study in Table I.
@@ III Dataset
TABLE I: Summary statistics of the dataset used in the analysis. Retweets have likes, replies, and quotes counts equal to zero, and they inherit retweet and view counts from the original tweets. Therefore, they have been used only to compute ideology in the Ukraine-Russia War dataset and excluded from the rest of the analysis.
IV Methods
@@ IV-A Pre-Processing
IV-B Factuality and Political Bias Estimation
To test the dependence of visibility variations on content-level features like political bias and factuality (cf. RQ1 in Section I), we rely on third-party classifications to infer the political bias and factuality of the news sources referenced in the debate. More precisely, we use a dataset obtained from Media Bias/Fact Check (MBFC), an independent fact-checking organization that categorizes news outlets based on their factuality and political bias. [^4] https://mediabiasfactcheck.com/. This includes 2,190 different news outlets, along with their domain names, political leaning, and factuality, and classifies outlets across the political spectrum, from “extreme left” to “extreme right.” Additionally, some outlets are labeled as “questionable” or “conspiracy-pseudoscience” if they frequently publish misinformation, false content, or endorse conspiracy theories.
@@ IV-C Influencers Selection
IV-D Users’ Ideological Stance Estimation
@@
IV-E Claim Detection and Tracking
@@
IV-F P-score
$\small\text{p-score}=\frac{\text{view count}}{\text{number of followers}}$ (1)
V Detecting Visibility Variations
In this section, we set out to answer our three research questions (cf. Section I) related to studying the occurrence of shadow banning at content, user, and network levels.
@@ V-A RQ1: Content Level
We classify tweets based on the domain they refer to using the MBFC dataset (see Section IV-B), dividing them into five categories: “News Outlets,” “No Domain,” “Other,” “Other Social,” and “Twitter.” These refer, respectively, to tweets that contain a link to a news outlet, no link, a link to an unclassified website, a link to another social platform, and quoting a tweet.
@@
TABLE II: Tweets by content type for distributions from Figure 2.
In Figure 2, we present the distribution of p-scores across tweet categories for both datasets (also see Table II). In both cases, the median p-score (marked by the red line) for the “No Domain” category is higher than for all others, confirming observations from Figure 1. Specifically, tweets from news outlets exhibit the lowest median p-scores (0.033 for Ukraine-Russia, 0.012 for US Elections), while “No Domain” content shows significantly higher medians (Ukraine-Russia: 0.246; US Elections: 0.056). This further supports the hypothesis that Twitter’s algorithm favors content that does not contain external URLs. Interestingly, tweets linking to unclassified websites have median p-scores close to those of news outlets (Ukraine-Russia: 0.036; US Elections: 0.014), whereas tweets quoting other tweets or linking to other social platforms are slightly less penalized (median p-scores of 0.097 and 0.150 for Ukraine-Russia and 0.029 and 0.024 for US Elections). To validate these observations, we perform a Mann-Whitney U test between all pairs of content categories to assess whether the values in one group tend to be greater than those in another [30]. The Mann-Whitney U test does not assume normality and instead evaluates whether there is a statistically significant shift between distributions. In our case, a low p-value indicates strong evidence that the values from one group are systematically higher than those from another. For both the Ukraine-Russia and US Elections datasets, the tests confirm that the “No Domain” category statistically significantly dominates all others (p-value <2.2e16<2.2e-16 for all comparisons).
@@
TABLE III: Number of tweets by political bias and factuality for the distributions from Figure 3.
Political Bias and Factuality. Next, we analyze the dependence of visibility on the political bias and factuality of the content being shared. We apply the same methodology to assess whether visibility varies based on the political bias and factuality of the sources referenced in tweets, for both the Ukraine-Russia War (Figure 3(a)3(b)) and the 2024 US Elections (Figure 3(c)3(d)) debates (also see Table III). For both datasets, the results show a relatively contained variation in visibility across both political bias (Ukraine-Russia: min 0.019, max 0.05; US Elections: min 0.0081, max 0.015) and factuality (Ukraine-Russia: min 0.018, max 0.04; US Elections: min 0.0086, max 0.024), suggesting that the algorithm does not systematically penalize content based on these attributes. This is also confirmed via the Mann-Whitney U test, which did not find a dominant category for political leaning in the Ukraine-Russia dataset (all categories exhibit p-values >> 0.05 in at least one comparison). As for factuality, the tests highlight that the “Mixed” and “Low” categories tend to be penalized (maximum p-values <2.2e16<2.2e-16 and 0.0019, respectively), but reject the hypothesis that “Very Low” content is consistently outperformed by “High” or “Very High” categories. Similarly, in the US Elections dataset, there is no evidence of clear dominance among political leanings (all categories have p-values >> 0.05 in at least one comparison), although the tests detect the lower performance of the “Left” category (p-value <2.2e16<2.2e-16). Interestingly, for factuality, “Very High” content outperforms all other categories (p-values <2.2e16<2.2e-16 in all comparisons), while no clear dominance emerges among the remaining classes. However, the number of tweets in the “Very High” category is considerably smaller than in others (see Table III), suggesting an occasional rather than systematic consumption of highly factual content in the US Elections debate.
Topics. Next, we use the textual content of tweets to investigate whether the different topics being discussed experienced varying levels of visibility. Following the claim detection and tracking methodology discussed in Section IV-E, for the Ukraine-Russia dataset, we first filter out 2,735 seed tweets making a claim. These seed tweets are further categorized into five different topics: 1) Europe/NATO Military Support to Ukraine, 2) Criticism of US Military Support to Ukraine, 3) US Military Support to Ukraine, 4) War Related Reporting and Events, and 5) Pro-Russian Support. Keyword extraction and the subsequent matching procedure return 151,697 tweets discussing claims related to the seed topics. We applied the same procedure for US elections dataset, categorizing the seeds into the following topics: 1) Anti-Trump Campaign Messaging, 2) International Conflicts & Electoral Impact, 3) Policy Debates & Issues, 4) Pro-Trump Campaign Messaging, 5) Public Reaction & Media Coverage, 6) Scandals, Conspiracies & Legal Controversies
@@
V-B RQ2: User Level
Inferring Users’ Ideology. To perform account-level analysis, we first need to infer users’ attributes, such as their stance on the debated issue. To this end, we use latent ideology estimation (see Section IV-D) to classify users’ ideological stances. Given a set of prominent accounts known as influencers (see Section IV-C), this technique allows us to infer users’ stances based on interaction information. Matrix decomposition maps users who interacted with a similar set of influencers close to each other in a real space, while users with fewer common influencers are positioned farther apart.
Due to the absence of retweet statistics for the 2024 US Elections dataset, we estimate latent ideology using retweet interactions for the Ukraine-Russia and reply interactions for the US Elections datasets. As mentioned, we do so as the latter only includes a few hundred retweets that did not have a valid view count. Although users’ estimaton results a bit noisier, the use of replies does not affect the overall ability to infer accounts’ stances as shown in Figure 6 and Figure 7, where we can still distinguish between influencers pro/against aid and pro/against Trump. Both datasets exhibit two communities positioned at opposite ends of the [1,1][-1,1] space, revealing polarized debates with bimodal opinion distributions, as also highlighted by Baqir et al. [5] in the context of the Ukraine-Russia war and by the Dip test as reported in Section IV-C.
@@ V-B RQ2: User Level
Figure 7: Distribution of p-score for influencers according to their popularity (i.e., number of followers). Only influencers with more than five original tweets were considered, and accounts with less than 50,000 followers were anonymized.
(a) Ukraine-Russia War
(b) 2024 US Elections
@@
Influences p-score distributions. Although we have ruled out the penalization of specific ideologies, visibility alterations may still occur for specific accounts. This is suggested by the presence of outliers among influencers in Figure 8, which may indicate accounts experiencing either reduced or increased content circulation due to algorithmic visibility adjustments. To investigate the presence of algorithmic interventions and to mitigate the influence of potential confounding factors, we exclusively focus on influencers and divide them based on their popularity in different tiers, as done in recent work [40]. Figure 7 displays the p-score distribution for the accounts used as influencers in the analysis of the two debates that posted more than five original pieces of content, with accounts grouped according to the popularity tier they belong to.
@@
Figure 9: P-score versus retweets per impression count for content produced by two news agencies with opposing ideological stances in the Ukraine-Russia war debate (a) and by the two presidential candidates in the US Election discussion (b). The volume of content produced by the four accounts are 553(KyivIndependent), 258(RT_com), 223(Donald J. Trump) and (467)Kamala Harris.
@@ VII Conclusion
Acknowledgments. A.G. and G.S. gratefully acknowledge support from, respectively, CY4GATE and NSF Grant CNS-2419829.

### v4

# Alessandro Galeazzi1, Pujan Paudel2, Mauro Conti1,3, Emiliano De Cristofaro4, Gianluca Stringhini2

@@ 1 Introduction
Beyond anecdotal evidence, researchers have begun looking into shadow banning through technical and statistical analyses (see Section 2). Overall, prior work suggests that shadow banning is relatively widespread across social media platforms [29, 7]. However, the lack of transparency in applying these techniques raises serious concerns. This prompts the need for systematic techniques to detect and measure visibility alterations and assess whether platforms employ shadow banning to favor specific narratives, limit dissenting voices, or prioritize corporate interests.
@@
In this paper, we aim to address this gap and evaluate the presence of visibility alterations in the wild using view counts, a largely underexplored metric in related work introduced on Twitter in 2022 and once available through the now mostly defunct API. As academic access to the API has been restricted since June 2023 [34, 22], we turn to two publicly available datasets that include view counts: 1) almost 35M tweets related to the 2024 US Presidential Election, released by Balasubramanian et al. [4], and 2) over 17M tweets related to military aid in the Ukraine-Russia war, released by Baqir et al. [5]. [^1] As Twitter rebranded to X in July 2023, tweets are now called posts. Since one of our datasets was collected before the rebranding and one after, to ease presentation, in the rest of the paper, we stick to “tweets.” Our methodology relies on data analysis and network science techniques to uncover traces of shadow banning at the content, user, and network levels, shedding light on visibility alterations. Although the inaccessibility of the platform’s recommendation algorithm prevents us from establishing an explicit causal link to algorithmic design, our findings provide reasonable evidence in support of this hypothesis. While alternative explanations are possible, we discuss why they are unlikely to hold in Section 6.
@@ 2 Related Work
Specific to shadow banning, Le Merrer et al. [29] investigate Twitter algorithmic interventions to reduce the visibility of certain content or accounts without outright removal. They highlight widespread traces of visibility interventions across various user groups and communities. Compared to our work, their study focuses on account-level interventions based on appearances in suggestions, search results, or discussion cascades, without considering alterations at the content level.
@@
Table 1: Summary statistics of the dataset used in the analysis. Retweets have likes, replies, and quotes counts equal to zero, and they inherit retweet and view counts from the original tweets. Therefore, they have been used only to compute ideology in the Ukraine-Russia War dataset and excluded from the rest of the analysis.
3 Dataset
2024 US Elections. The second dataset was collected by Balasubramanian et al. [4] using a custom-built scraper. [^3] The dataset is available at https://github.com/sinking8/usc-x-24-us-election. The authors curated a list of 44 keywords related to the 2024 US Elections, including ‘2024 Elections,’ ‘Donald Trump,’ ‘letsgobrandon,’ ‘ultramaga,’ etc. Following retrieval, we filter the data as discussed in Section 4.1 and obtain approximately 35 million tweets with valid view counts by 3.8 million users.
As we apply several transformations to the datasets throughout the analysis (see Section 4), we provide statistics for the datasets used in the main steps of this study in Table 1.
4 Methods
@@ 4.1 Pre-Processing
4.2 Factuality and Political Bias Estimation
To test the dependence of visibility variations on content-level features like political bias and factuality (cf. RQ1 in Section 1), we rely on third-party classifications to infer the political bias and factuality of the news sources referenced in the debate. More precisely, we use a dataset obtained from Media Bias/Fact Check (MBFC), an independent fact-checking organization that categorizes news outlets based on their factuality and political bias. [^4] https://mediabiasfactcheck.com/. This includes 2,190 different news outlets, along with their domain names, political leaning, and factuality, and classifies outlets across the political spectrum, from “extreme left” to “extreme right.” Additionally, some outlets are labeled as “questionable” or “conspiracy-pseudoscience” if they frequently publish misinformation, false content, or endorse conspiracy theories.
@@ 4.3 Influencers Selection
4.4 Users’ Ideological Stance Estimation
@@
4.5 Claim Detection and Tracking
@@
4.6 P-score
$\text{p-score}=\frac{\text{view count}}{\text{number of followers}}$ (1)
5 Detecting Visibility Variations
In this section, we set out to answer our three research questions (see Section 1) related to studying the occurrence of shadow banning at content, user, and network levels.
@@ 5.1 RQ1: Content Level
We classify tweets based on the domain they refer to using the MBFC dataset (see Section 4.2), dividing them into five categories: “News Outlets,” “No Domain,” “Other,” “Other Social,” and “Twitter.” These refer, respectively, to tweets that contain a link to a news outlet, no link, a link to an unclassified website, a link to another social platform, and a link to other Twitter content.
@@
In Figure 2, we present the distribution of p-scores across tweet categories for both datasets (also see Table 2). In both cases, the median p-score (marked by the red line) for the “No Domain” category is higher than for all others, confirming observations from Figure 1. Specifically, tweets from news outlets exhibit the lowest median p-scores (0.033 for Ukraine-Russia, 0.012 for US Elections), while “No Domain” content shows significantly higher medians (Ukraine-Russia: 0.246; US Elections: 0.056). This further supports the hypothesis that Twitter’s algorithm favors content that does not contain external URLs. Interestingly, tweets linking to unclassified websites have median p-scores close to those of news outlets (Ukraine-Russia: 0.036; US Elections: 0.014), whereas tweets quoting other tweets or linking to other social platforms are slightly less penalized (median p-scores of 0.097 and 0.150 for Ukraine-Russia and 0.029 and 0.024 for US Elections). To validate these observations, we perform a Mann-Whitney U test between all pairs of content categories to assess whether the values in one group tend to be greater than those in another [30]. The Mann-Whitney U test does not assume normality and instead evaluates whether there is a statistically significant shift between distributions. In our case, a low p-value indicates strong evidence that the values from one group are systematically higher than those from another. For both the Ukraine-Russia and US Elections datasets, the tests confirm that the “No Domain” category statistically significantly dominates all others (p-value <2.2e16<2.2e-16 for all comparisons).
@@
Table 2: Tweets by content type for distributions from Figure 2.
@@
Table 3: Number of tweets by political bias and factuality for the distributions from Figure 3.
@@
Political Bias and Factuality. Next, we analyze the dependence of visibility on the political bias and factuality of the content being shared. We apply the same methodology to assess whether visibility varies based on the political bias and factuality of the sources referenced in tweets, for both the Ukraine-Russia War (Figure 3(a)3(b)) and the 2024 US Elections (Figure 3(c)3(d)) debates (also see Table 3). For both datasets, the results show a relatively contained variation in visibility across both political bias (Ukraine-Russia: min 0.019, max 0.05; US Elections: min 0.0081, max 0.015) and factuality (Ukraine-Russia: min 0.018, max 0.04; US Elections: min 0.0086, max 0.024), suggesting that the algorithm does not systematically penalize content based on these attributes. This is also confirmed via the Mann-Whitney U test, which did not find a dominant category for political leaning in the Ukraine-Russia dataset (all categories exhibit p-values >> 0.05 in at least one comparison). As for factuality, the tests highlight that the “Mixed” and “Low” categories tend to be penalized (maximum p-values <2.2e16<2.2e-16 and 0.0019, respectively), but reject the hypothesis that “Very Low” content is consistently outperformed by “High” or “Very High” categories. Similarly, in the US Elections dataset, there is no evidence of clear dominance among political leanings (all categories have p-values >> 0.05 in at least one comparison), although the tests detect the lower performance of the “Left” category (p-value <2.2e16<2.2e-16). Interestingly, for factuality, “Very High” content outperforms all other categories (p-values <2.2e16<2.2e-16 in all comparisons), while no clear dominance emerges among the remaining classes. However, the number of tweets in the “Very High” category is considerably smaller than in others (see Table 3), suggesting an occasional rather than systematic consumption of highly factual content in the US Elections debate.
@@
Topics. Next, we use the textual content of tweets to investigate whether the different topics being discussed experienced varying levels of visibility. Following the claim detection and tracking methodology discussed in Section 4.5, for the Ukraine-Russia dataset, we first filter out 2,735 seed tweets making a claim. These seed tweets are further categorized into five different topics: 1) Europe/NATO Military Support to Ukraine, 2) Criticism of US Military Support to Ukraine, 3) US Military Support to Ukraine, 4) War Related Reporting and Events, and 5) Pro-Russian Support. Keyword extraction and the subsequent matching procedure return 151,697 tweets discussing claims related to the seed topics. We applied the same procedure for US elections dataset, categorizing the seeds into the following topics: 1) Anti-Trump Campaign Messaging, 2) International Conflicts & Electoral Impact, 3) Policy Debates & Issues, 4) Pro-Trump Campaign Messaging, 5) Public Reaction & Media Coverage, 6) Scandals, Conspiracies & Legal Controversies
@@
5.2 RQ2: User Level
Inferring Users’ Ideology. To perform account-level analysis, we first need to infer users’ attributes, such as their stance on the debated issue. To this end, we use latent ideology estimation (see Section 4.4) to classify users’ ideological stances. Given a set of prominent accounts known as influencers (see Section 4.3), this technique allows us to infer users’ stances based on interaction information. Matrix decomposition maps users who interacted with a similar set of influencers close to each other in a real space, while users with fewer common influencers are positioned farther apart.
Due to the absence of retweet statistics for the 2024 US Elections dataset, we estimate latent ideology using retweet interactions for the Ukraine-Russia and reply interactions for the US Elections datasets. As mentioned, we do so as the latter only includes a few hundred retweets that did not have a valid view count. Although users’ estimaton results a bit noisier, the use of replies does not affect the overall ability to infer accounts’ stances. Indeed, the bar plots in Figure 6 show the distribution of ideology scores for users and influencers in the Ukraine–Russia War (panel a) and the US Elections (panel b), together with the distribution of ideology scores for repliers/retweeters of a subset of ten influencers in each debate. Both datasets exhibit two communities positioned at opposite ends of the [1,1][-1,1] space, revealing polarized debates with bimodal opinion distributions, as also highlighted by Baqir et al. [5] in the context of the Ukraine-Russia war and by the Dip test as reported in Section 4.3.
@@ 5.2 RQ2: User Level
Influencers p-score distributions. Although we have ruled out the penalization of specific ideologies, visibility alterations may still occur for specific accounts. This is suggested by the presence of outliers among influencers in Figure 7, which may indicate accounts experiencing either reduced or increased content circulation due to algorithmic visibility adjustments. To investigate the presence of algorithmic interventions and to mitigate the influence of potential confounding factors, we exclusively focus on influencers and divide them based on their popularity in different tiers, as done in recent work [40]. Figure 8 displays the p-score distribution for the accounts used as influencers in the analysis of the two debates that posted more than five original pieces of content, with accounts grouped according to the popularity tier they belong to.
@@
Figure 8: Distribution of p-score for influencers according to their popularity (i.e., number of followers). Only influencers with more than five original tweets were considered, and accounts with fewer than 50,000 followers were anonymized.
(a) Ukraine-Russia War
(b) 2024 US Elections
Figure 9: P-score vs. retweets per impression count for content by two news agencies with opposing ideological stances. The volume of content produced by the four accounts is 553 (KyivIndependent), 258 (RT_com), 223 (Donald J. Trump), and 467 (Kamala Harris).
@@ 7 Conclusion
Acknowledgments. A.G. gratefully acknowledges the financial support provided by CY4GATE, by the National Recovery and Resilience Plan (NRRP) project “Securing Software Platforms – SOP” (CUP H73C22000890001), and by the European Union under the NRRP, Mission 4 Component 2 Investment 1.3 – Call for Proposals No. 341 of March 15, 2022, Italian Ministry of University and Research – NextGenerationEU, Project Code PE00000014, Concession Decree No. 1556 of October 11, 2022 (CUP D43C22003050001), “SEcurity and RIghts in the CyberSpace (SERICS) – Spoke 2 Misinformation and Fakes: DEcision supporT systEm foR cybeR intelligENCE (Deterrence). G.S.’s work was partially supported by NSF Grant CNS-2419829.